A machine that manufactures automobile parts produces defective parts
13% of the time. If 7 parts produced by this machine are randomly selected, what is the probability that fewer than 3 of the parts are defective?

Answer by Boreal(15235)

This means 0, 1, or 2 parts are defective, and there is an 87% probability of not-defective. Look at the likelihood of getting no defectives.
for 0, the probability is 0.87^7=0.3773
for 1 it is 7*0.87^6*0.13=0.3945
for 2 it is 21*0.87^5*0.13^2=0.1769
that sum is 0.9487 and is the answer.